OpenStudy (anonymous):

Find the product of the sum and difference of two terms: (4+5y^3)(4-5y^3) My answer of 16+25y^3 is wrong.

OpenStudy (anonymous):

(+/-)80 y^3. let 4 = a, 5y^3 = b. sum = a+b+a-b = 2a. difference is => (a+b)-(a-b) = 2b or (a-b)-(a+b) = -2b product is 4ab or -4ab => 80y^3 or -80y^3